If the focus and vertex of a parabola arr thr points (0,2) and (0,4) respectively, then its equation is

Grade:12

1 Answers

Arun
6 years ago
The coordinates of the vertex is (0,4)
The coordinates of the focus is (0,2)
It is clear that the vertex and the focus lies on the positive side of the y-axis.
Hence the curve is open downwards.
The equation of the form (x-h)^2=4a(y-k)
(ie) 
(x-0)^2 = -4*2*(y-4)
On simplifying we get,
x^2 = -8y +32
x^2 + 8y = 32